# Oscillatons in Scalar-Field Dark Matter from a Full Fourier Expansion of an Exponential Potential

### v2

@@ III.6 Energy conditions and the interpretation of negative pressures
Then the components of stress eergy tensor up to any desired order of expansion, in this text up to order four/six for metric prefactors, $e^{-(\nu\pm\mu)}$/ potential, $e^{-2\lambda\sigma}$, respectively can be obtained as:
@@
Figure 7: Period-averaged conterparts of energy densiity/flux and Radia/tangential pressures for $\sigma(0)=\frac{1.2}{\sqrt{2}}$, $\lambda=1$. Left: $\langle\,\rho_{r}(t,x)\rangle_{t}=\langle\tfrac{1}{2}\!\big[e^{-(\nu-\mu)}\dot{\Phi}^{2}+e^{-(\nu+\mu)}\Phi^{\prime 2}+2V(\Phi)\big]\rangle_{t}$$\langle\,P_{r}(t,x)\rangle_{t}=\langle\tfrac{1}{2}\!\big[e^{-(\nu-\mu)}\dot{\Phi}^{2}+e^{-(\nu+\mu)}\Phi^{\prime 2}-2V(\Phi)\big]\rangle_{t}$ and $\langle\,P_{\perp}(t,x)\rangle_{t}=\langle\tfrac{1}{2}\!\big[e^{-(\nu-\mu)}\dot{\Phi}^{2}-e^{-(\nu+\mu)}\Phi^{\prime 2}-2V(\Phi)\big]\rangle_{t}$. Right: $\langle j_{r}(t,x)\rangle_{t}=\langle\,e^{-(\nu+\mu)}\dot{\Phi}\Phi^{\prime}\rangle_{t}$. The GR interpretation of negative parts of quantities can be discussed in Sec. III.6 via WEC/NEC/SEC diagnostics. The average energy flux for dominate frequencies $2\omega\,t-8\omega\,t$, is zero at fixed x, indicating that there is no net mass transfer in a stationary oscillaton.
In a local orthonormal tetrad adapted to $(t,r,\theta,\varphi)$ we use the standard pointwise inequalities (see for exxample (50, ) and references therein)

### v3

@@ III.6 Energy conditions and the interpretation of negative pressures
Then the components of stress energy tensor up to any desired order of expansion, in this text up to order four/six for $e^{-(\nu\pm\mu)}$$e^{-2\lambda\sigma}$, respectively can be obtained as:
@@
Figure 7: Period-averaged counterparts of energy densiity/flux and Radia/tangential pressures for $\sigma(0)=\frac{1.2}{\sqrt{2}}$, $\lambda=1$. Left: $\langle\,\rho_{r}(t,x)\rangle_{t}=\langle\tfrac{1}{2}\!\big[e^{-(\nu-\mu)}\dot{\Phi}^{2}+e^{-(\nu+\mu)}\Phi^{\prime 2}+2V(\Phi)\big]\rangle_{t}$$\langle\,P_{r}(t,x)\rangle_{t}=\langle\tfrac{1}{2}\!\big[e^{-(\nu-\mu)}\dot{\Phi}^{2}+e^{-(\nu+\mu)}\Phi^{\prime 2}-2V(\Phi)\big]\rangle_{t}$ and $\langle\,P_{\perp}(t,x)\rangle_{t}=\langle\tfrac{1}{2}\!\big[e^{-(\nu-\mu)}\dot{\Phi}^{2}-e^{-(\nu+\mu)}\Phi^{\prime 2}-2V(\Phi)\big]\rangle_{t}$. Right: $\langle j_{r}(t,x)\rangle_{t}=\langle\,e^{-(\nu+\mu)}\dot{\Phi}\Phi^{\prime}\rangle_{t}$. The GR interpretation of negative parts of quantities can be discussed in Sec. III.6 via WEC/NEC/SEC diagnostics. The average energy flux for dominate even frequencies between $2\omega\,t$ and $8\omega\,t$, is zero at fixed x, indicating that there is no net mass transfer in a stationary oscillaton.
In a local orthonormal tetrad adapted to $(t,r,\theta,\varphi)$ we use the standard pointwise inequalities (see for example (50, ) and references therein)
@@
Finally the mass seen by an observer at infinity obtained straightly through:
$M_{\Phi}=\dfrac{1}{8\pi}m_{\Phi}^{2}M_{Pl}^{2}\int_{0}^{\infty}\langle\rho(t,x)\rangle_{t}\,4\pi\,x^{2}\,dx,$ (38)
shows that the mass and radius of the oscillaton are inversely proportional to $\sigma_{c}(0)$, meaning that smaller values of radial amplitude lead to more massive oscillaton with larger radius and vice vera (see the following table 1).
$\sigma_{c}(0)$ $\frac{1.2}{\sqrt{2}}$     $\frac{1.3}{\sqrt{2}}$     $\frac{1.4}{\sqrt{2}}$     $\frac{1.5}{\sqrt{2}}$
$M_{\Phi}(\frac{1}{2}m_{\Phi}^{2}\,M_{Pl}^{2})$ 89.08     51.00     42.18     29.54
x(radius) 1.31     1.23    1.13     0.999
Table 1: Oscillaton mass and its radius for different values of $\sigma_{c}(0)$ as radial amplitude of scalar field.